# Effect of Synthetic Jets Actuator Parameters on Deep Reinforcement Learning-Based Flow Control Performance in a Square Cylinder

### v2

@@ Abstract
We utilize deep reinforcement learning (DRL) algorithms to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of a square cylinder for active flow control. Through DRL-based active flow control (AFC) technology, we significantly reduce the lift and drag coefficients of the square cylinder at Reynolds number ($Re$) $=100$ and $Re=500$, while completely suppressing vortex shedding in the wake flow field. Additionally, we conduct a sensitivity analysis of the position and width parameters of the synthetic jets regarding flow control performance. Our observations indicate that positioning the synthetic jets near the trailing edge corners of the square cylinder, rather than the leading edge corners, can completely suppress vortex shedding, resulting in more stable lift and drag coefficients in the controlled flow. When the synthetic jets are positioned at the trailing edge corners, flow control reduces the mean drag coefficient by 14.4% and the standard deviation of the lift coefficient by 86.1% for the baseline flow at $Re=100$. For the baseline flow at $Re=500$, flow control reduces the mean drag coefficient by 51.4% and the standard deviation of the lift coefficient by 90.5%. At both Reynolds numbers, vortex shedding in the wake flow field is completely suppressed. Furthermore, using narrower synthetic jets results in a lower reduction rate of the standard deviations of the lift and drag coefficients, while increasing the mean and standard deviation of the mass flow rate of the jets used for flow control. This study provides guidance on optimizing the width and position of synthetic jets for DRL-based active flow control.
@@ I INTRODUCTION
In the backdrop of advancements in sophisticated algorithms, computational hardware, open-source software and the integration of vast amounts of data in the era of big data, scientific inquiry is transitioning from first principles to data-driven methods.Brunton, Noack, and Koumoutsakos (2020) The field of fluid dynamics is also benefiting from this trend, with machine learning offering sophisticated algorithms capable of handling high-dimensional, large-scale fluid dynamics data, establishing mathematical modeling frameworks, and providing integrated, modular packages and libraries as open-source resources. Machine learning excels in tasks like feature extraction, modal analysis, and data mining, providing techniques for extracting insights from datasets.Mahesh (2020) Experimental fluid dynamics measurements and numerical simulation generate extensive datasets, making them ideal candidates for techniques capable of handling high-dimensional, large-scale data.Manohar et al. (2018) Furthermore, machine learning algorithms can embed physical information from the field of fluid dynamics and adaptively perform tasks related to execution, control, and optimization.Perdikaris, Venturi, and Karniadakis (2016); Bewley, Moin, and Temam (2001) The field of flow control is rapidly advancing, propelled by the development of state-of-the-art machine learning algorithms and unprecedented volumes of data from high-precision numerical simulations across multiple spatiotemporal scales.Brunton, Noack, and Koumoutsakos (2020); Manohar et al. (2018); Perdikaris, Venturi, and Karniadakis (2016); Bewley, Moin, and Temam (2001)
@@
Active Flow Control (AFC) is an advanced fluid management strategy aimed at enhancing fluid properties and performance through proactive interventions in the flow field. By introducing energy or momentum, AFC alters the fluid’s natural state, employing techniques such as fluid injection and suction, synthetic jets, and electromagnetic control.Cattafesta and Sheplak (2011); Aram et al. (2018) Synthetic jet technology produces jets by cyclically inhaling and expelling fluid from a fixed location, requiring no external energy source.Glezer and Amitay (2002) This method offers several advantages, including independence from external fluid sources, high controllability, low energy consumption, and suitability for operation in narrow or complex spaces.Jahanmiri (2010); Smith and Glezer (1998) These advantages render synthetic jets particularly promising tools for present and future flow control technologies.François-Lavet et al. (2018); Xie et al. (2023) AFC is widely applied across various sectors, including aerospace, automotive, energy production, and environmental engineering, where it serves to reduce drag, control vortices, enhance lift, and stabilize flows.Scott Collis et al. (2004); Smith and Swift (2003) Despite its considerable advantages, the implementation and practical application of AFC face challenges due to the system’s strong nonlinearity, high dimensionality, and time delays. Recent advancements in machine learning algorithms and control theory are propelling progress in adaptive, real-time intelligent control, facilitating more effective management of complex flow dynamics and further augmenting the capabilities of AFC technologies.Brunton and Noack (2015); Luchtenburg (2021); Brunton, Proctor, and Kutz (2016)
@@
Deep Reinforcement Learning (DRL) merges the capabilities of Deep Learning (DL) for processing high-dimensional data with the decision-making and policy optimization strengths of Reinforcement Learning (RL).Mnih et al. (2013); LeCun, Bengio, and Hinton (2015) DRL is recognized for its ability to learn and adapt autonomously, enabling end-to-end learning that is highly applicable across complex environments. DRL has demonstrated exceptional performance in a wide range of domains, including game AI, robotic control, autonomous driving, resource management and scheduling, and financial trading. It is capable of autonomously learning optimal decision-making strategies in complex environments, surpassing human-level capabilities.Mnih et al. (2015); Lillicrap et al. (2019); Henderson et al. (2019) DRL is a sophisticated blend of DL and RL, two powerful branches of machine learning. It combines the feature extraction capabilities of DL for complex high-dimensional data with the decision-making abilities of RL through interactions with the environment.Li (2018); Kaiser et al. (2024) This synergy allows DRL to excel in environments that require both advanced perception and precise control. Despite challenges related to sample efficiency and stability, DRL demonstrates remarkable potential on a variety of challenging decision-making problems.Arulkumaran et al. (2017); Kaiser et al. (2024)
@@
Given the significant potential of DRL in control and optimization, a large number of scholars have been applying it in the field of flow control. Rabault et al.Rabault et al. (2019) pioneered the application of DRL to AFC technologies, achieving an 8% reduction in drag around a cylinder at $Re=100$. This seminal work sparked interest in applying DRL techniques within the field of fluid dynamics. Tang et al.Tang et al. (2020a) focused on the robustness across different Reynolds numbers, while Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021) explored the impact of probe distribution on control performance. Li and ZhangLi and Zhang (2022) applied DRL-based AFC to cylinders under various blockage ratios, integrating flow physics mechanisms during the control process to achieve significant outcomes. Fan et al.Fan et al. (2020) developed an active flow control strategy based on DRL through experimental fluid dynamics, demonstrating the effectiveness of applying reinforcement learning in experimental fluid mechanics. Rabault and KuhnleRabault and Kuhnle (2019a) and Jia and XuJia and Xu (2024a) addressed issues related to parallel strategies and training when coupling Computational Fluid Dynamics (CFD) codes with DRL framework in parallel computing environments. Wang et al.Wang et al. (2022) developed an open source library that integrated CFD code OpenFoam with DRL framework TensorForce. DRL has demonstrated excellent control performance in the field of AFC, sparking increasing exploration. There has been applications from low to relatively high Reynolds numbers, transitions from 2D to 3D configurations, shifts from numerical simulations to fluid experiments, and geometric changes in the body of interest from circular to square bodies. While there is a plethora of related work, we will not delve into each individually. Instead, we summarize in Table. 1 some relevant research on the application of DRL-based AFC methods in the context of flow around bluff bodies.
@@
100 Cylinder Rabault et al.Rabault et al. (2019) Synthetic Jets PPO FeniCS 8% -
@@
100 Cylinder Castellanos et al.Castellanos et al. (2022) Synthetic Jets LGPC/PPO FeniCS 8% -
100 Cylinder Jia and XuJia and Xu (2024b) Synthetic Jets PPO OpenFOAM 8% YES
100 Cylinder Ren, Wang, and TangRen, Wang, and Tang (2021) WSLB PPO LBM - -
120 Cylinder Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021) Synthetic Jets S-PPO-CMA FastS 18.4% -
@@
100-300 Cylinder He et al.He et al. (2023) Synthetic Jets PPO OpenFOAM 6%-24% -
100-400 Cylinder Tang et al.Tang et al. (2020b) Synthetic Jets PPO FeniCS 6%-39% -
@@
10,160 Cylinder Fan et al.Fan et al. (2020) Cylinders TD3 Lilypad 30% -
100 Ellipse Jia and XuJia and Xu (2024b) Synthetic Jets PPO OpenFOAM 16% YES
@@
100 Square Xia et al.Xia et al. (2024) Synthetic Jets SAC and TQC FEniCS 17.3% -
100-400 Square Jia and XuJia and Xu (2024c) Synthetic Jets SAC OpenFOAM 14%-47% YES
@@
The research in the above table demonstrate the wide range of applications of DRL in the field of active flow control. These research significantly extend beyond the initial study of flow past a cylinder at $Re=100$ by Rabault et al.Rabault et al. (2019). However, the effectiveness of DRL control algorithm varies across different work. This is especially evident among the body of research on square cylinders. As the sharp corners and straight edges force the fluid to separate at the leading edge, flows around square cylinders are characterized by complex wake structures and shedding vortices.Bai and Alam (2018a); Yoon, Yang, and Choi (2010); Sen, Mittal, and Biswas (2011) The highly nonlinear and unique flow features makes it extremely challenging to determine the optimal control strategy. Hence, many different control experiments have been devised heuristically, leading to varying degrees of control effectiveness.
Wang et al.Wang et al. (2022) initially applied the Soft Actor-Critic (SAC) algorithm to active flow control of a square cylinder, achieving a 13.7% reduction in drag at $Re=100$. Despite this significant reduction, vortex shedding behind the cylinder is still evident. Xia et al.Xia et al. (2024) conduct active flow control around a square cylinder at $Re=100$ using RL, training with an augmented state that includes both current and past measurements and actions, which are represented by a nonlinear auto-regressive exogenous model to address issues of partial observability. Jia and XuJia and Xu (2024c) employed the SAC algorithm to regulate the flow from synthetic jets positioned on the upper and lower surfaces of a constrained square cylinder at $Re=100$, 200, 300, and 400. The research results demonstrated that the active control strategy based on the SAC algorithm effectively reduces drag by approximately 14.4%, 26.4%, 38.9%, and 47.0% respectively. A distinctive feature of this study compared to others is its success in suppressing vortex shedding behind the square cylinder across the range of Reynolds numbers from 100 to 400. This DRL-based AFC delivers an energy-efficient control strategy that not only reduces drag and suppresses lift but also effectively suppresses vortex shedding. Yan et al.Yan et al. (2024) investigated the impact of multiple jet actuators on drag and lift reduction for rectangular cylinders with four different aspect ratios at $Re=1,000$, employing DRL for control. Specifically, the deployment of four synthetic jets resulted in an average drag reduction of 63.2%, while eight jets achieved a 77.1% reduction. This study investigated various jet distribution strategies, emphasizing the importance of multi-jet actuation in enhancing aerodynamic performance.
Additionally, an intriguing research question that captures the interest of many scholars is the optimal placement of synthetic jets for control effectiveness. Particularly noteworthy are two studies that have arrived at markedly different conclusions, prompting further reflection and analysis within the research community.Chen et al. (2023); Yan et al. (2023) Chen et al.Chen et al. (2023) employ DRL-based AFC to mitigate vibrations induced by lift on a square cylinder at $Re=100$. Their study investigates the effect of jets positioning on control efficacy, examining positions near the leading edge, midpoints of the sidewalls, and near the trailing edge of the cylinder. Results indicate that for reducing lift-induced vibrations and drag coefficient, positioning the jets near the trailing edge is significantly superior to placement near the leading edge or midpoints of the sidewalls. Moreover, complete suppression of vortex shedding behind the square cylinder occurs only when the jets are placed near the trailing edge. Yan et al.Yan et al. (2023) examined the impact of symmetrically deploying jet actuators at both the front and trailing corners of a square cylinder under flow configurations at $Re=500$, 1,000 and 2,000. The findings reveal that positioning jets near the front corners provides more effective control over lift and drag coefficients compared to placement near the trailing corners. Despite these modifications, vortex shedding in the controlled flow was not suppressed at any of the three Reynolds numbers.
Given the current state of the art in DRL-based AFC of flow around square cylinders, it is evident that the there is a lack of parameterization for synthetic jets to reconcile the varying degrees of control effectiveness in literature. Consequently, our research is motivated by two main reasons. Firstly, there is currently no study that has successfully used AFC to achieve both efficient drag reduction and complete suppression of vortex shedding at Reynolds numbers greater than 500. Secondly, previous studies have shown that the effects of synthetic jets placement on control performance exhibit different outcomes at $Re=100$ and $Re=500$. We are interested in exploring whether it is possible, at $Re=100$ and $Re=500$, to use AFC based on DRL to achieve both efficient drag reduction and complete suppression of vortex shedding. Additionally, we design two synthetic jet placements and four synthetic jet widths to conduct a thorough analysis of their impacts on flow control performance. The parameterized design of control experiments and corresponding analysis aim to provide valuable references for the control of flow around square cylinders. The structure of this paper is as follows. In Sec. II, we introduce the DRL framework, which consists of a fluid simulation environment and the SAC agent. In Sec. III, we present and discuss the results. Initially, we describe the Baseline and the Base flow. Then, we discuss the convergence performance of DRL at $Re=100$ and $Re=500$. Finally, we analyze the effects of the placement and width of synthetic jets on control performance. In Sec. IV, we summarize the main conclusions of this study and discuss the broader implications of these findings.

### v3

@@ Abstract
We conduct an active flow control (AFC) study on the mass flow rate of synthetic jets on the upper and lower surfaces of a square cylinder using a deep reinforcement learning (DRL) algorithm, with a focus on investigating the influence of the position and width of the synthetic jets on the flow control performance. At Reynolds numbers ($Re$) of 100 and 500, it is found that our proposed method significantly reduced the lift and drag coefficients of the square cylinder, and completely suppressed vortex shedding in the wake. In particular, at $Re=100$, placing the synthetic jets near the tail corner was beneficial for reducing drag, with a maximum drag reduction rate of 14.4%. When $Re=500$, positioning the synthetic jets near the leading edge corner resulted in a maximum optimal drag reduction effect of 65.5%. This indicates that locating the synthetic jet at the main flow separation point can achieve optimal control. Furthermore, we notice that when the synthetic jets are positioned near the tail corner, vortex shedding can be completely suppressed. Additionally, a narrower width of the synthetic jets can enhance flow instability and increase the cost of flow control.
@@ I INTRODUCTION
In the backdrop of advancements in sophisticated algorithms, computational hardware, open-source software and the integration of vast amounts of data in the era of big data, scientific inquiry is transitioning from first principles to data-driven methods.DataDriven; Datadrivenscience The field of fluid dynamics is also benefiting from this trend, with machine learning offering sophisticated algorithms capable of handling high-dimensional, large-scale fluid dynamics data,Brunton, Noack, and Koumoutsakos (2020) establishing mathematical modeling frameworks, and providing integrated, modular packages and libraries as open-source resources.Brunton2015; Stevendynamical Machine learning excels in tasks like feature extraction, modal analysis, and data mining, providing techniques for extracting insights from datasets.Batta (2020); Datadrivenscience Experimental fluid dynamics measurements and numerical simulation generate extensive datasets, making them ideal candidates for techniques capable of handling high-dimensional, large-scale data.SCOTTCOLLIS2004237; DataDriven Furthermore, machine learning algorithms can embed physical information from the field of fluid dynamics and adaptively perform tasks related to execution, control, and optimization.Multifidelity; Bewley2001 The field of flow control is rapidly advancing, propelled by the development of state-of-the-art machine learning algorithms and unprecedented volumes of data from high-precision numerical simulations across multiple spatiotemporal scales.Brunton, Noack, and Koumoutsakos (2020); Multifidelity; Bewley2001
@@
Active flow control (AFC) is an advanced fluid management strategy aimed at enhancing fluid properties and performance through proactive interventions in the flow field. By introducing energy or momentum, AFC alters the fluid’s natural state, employing techniques such as fluid injection and suction, synthetic jets, and electromagnetic control.annurevCattafesta; Aram2018 Synthetic jet technology produces jets by cyclically inhaling and expelling fluid from a fixed location, requiring no external energy source.annurevfluidGlezer; SCOTTCOLLIS2004237 This method offers several advantages, including independence from external fluid sources, high controllability, low energy consumption, and suitability for operation in narrow or complex spaces.jahanmiri2010active; Smith1998 These advantages render synthetic jets particularly promising tools for present and future flow control technologies.MAL071; Aerospace2023 AFC is widely applied across various sectors, including aerospace, automotive, energy production, and environmental engineering, where it serves to reduce drag, control vortices, enhance lift, and stabilize flows.SCOTTCOLLIS2004237; smith2003comparison Despite its considerable advantages, the implementation and practical application of AFC face challenges due to the system’s strong nonlinearity, high dimensionality, and time delays. Recent advancements in machine learning algorithms and control theory are propelling progress in adaptive, real-time intelligent control, facilitating more effective management of complex flow dynamics and further augmenting the capabilities of AFC technologies.Brunton2015; Datadrivenscience; Stevendynamical
@@
Deep Reinforcement Learning (DRL) merges the capabilities of Deep Learning (DL) for processing high-dimensional data with the decision-making and policy optimization strengths of Reinforcement Learning (RL).mnih2013playing; lecun2015deep DRL is recognized for its ability to learn and adapt autonomously, enabling end-to-end learning that is highly applicable across complex environments. DRL has demonstrated exceptional performance in a wide range of domains, including game AI, robotic control, autonomous driving, resource management and scheduling, and financial trading. It is capable of autonomously learning optimal decision-making strategies in complex environments, surpassing human-level capabilities.mnih2015human; lillicrap2019continuous; henderson2019deep DRL is a sophisticated blend of DL and RL, two powerful branches of machine learning. It combines the feature extraction capabilities of DL for complex high-dimensional data with the decision-making abilities of RL through interactions with the environment.li2018deep; kaiser2024modelbased This synergy allows DRL to excel in environments that require both advanced perception and precise control. Despite challenges related to sample efficiency and stability, DRL demonstrates remarkable potential on a variety of challenging decision-making problems.Arulkumaran et al. (2017); kaiser2024modelbased
@@
Given the significant potential of DRL in control and optimization, a large number of scholars have been applying it in the field of flow control. rabault2019artificialrabault2019artificial pioneered the application of DRL to AFC technologies, achieving an 8% reduction in drag around a cylinder at $Re=100$. This seminal work sparked interest in applying DRL techniques within the field of fluid dynamics. tang2020robusttang2020robust focused on the robustness across different Reynolds numbers, while parisRobustFlowControl2021parisRobustFlowControl2021 explored the impact of probe distribution on control performance. Li and ZhangLi and Zhang (2022) applied DRL-based AFC to cylinders under various blockage ratios, integrating flow physics mechanisms during the control process to achieve significant outcomes. DixiapnasDixiapnas developed an active flow control strategy based on DRL through experimental fluid dynamics, demonstrating the effectiveness of applying reinforcement learning in experimental fluid mechanics. rabault2019acceleratingrabault2019accelerating and Jia and XuJia and Xu (2024) addressed issues related to parallel strategies and training when coupling Computational Fluid Dynamics (CFD) codes with DRL framework in parallel computing environments. Wang et al.Wang et al. (2022) developed an open source library that integrated CFD code OpenFoam with DRL framework TensorForce. DRL demonstrates excellent control performance in the field of AFC, sparking increasing exploration. There are applications from low to relatively high Reynolds numbers, transitions from 2D to 3D configurations, shifts from numerical simulations to fluid experiments, and geometric changes in the body of interest from circular to square bodies. While there is a plethora of related work, we will not delve into each individually. Instead, we summarize in Tab. 1 some relevant research on the application of DRL-based AFC methods in the context of flow around bluff bodies.
@@
100 Cylinder rabault2019artificialrabault2019artificial Synthetic Jets PPO FeniCS 8% -
@@
100 Cylinder castellanoscastellanos Synthetic Jets LGPC/PPO FeniCS 8% -
100 Cylinder jia2024deepjia2024deep Synthetic Jets PPO OpenFOAM 8% YES
100 Cylinder ren2021bluffren2021bluff WSLB PPO LBM - -
120 Cylinder parisRobustFlowControl2021parisRobustFlowControl2021 Synthetic Jets S-PPO-CMA FastS 18.4% -
@@
100-300 Cylinder he2023policyhe2023policy Synthetic Jets PPO OpenFOAM 6%-24% -
100-400 Cylinder tangRobusttangRobust Synthetic Jets PPO FeniCS 6%-39% -
@@
10,160 Cylinder DixiapnasDixiapnas Cylinders TD3 Lilypad 30% -
100 Ellipse jia2024deepjia2024deep Synthetic Jets PPO OpenFOAM 16% YES
@@
100 Square Xia2024Xia2024 Synthetic Jets SAC and TQC FEniCS 17.3% -
100-400 Square jia2024robustjia2024robust Synthetic Jets SAC OpenFOAM 14%-47% YES
@@
The flow around a square cylinder exhibits instabilities similar to those around a circular cylinder, but with different separation mechanisms and variations in lift, drag, and Strouhal number with Reynolds number. For square cylinders, the separation points are fixed at the leading or trailing edge, depending on the Reynolds number.Sohankar, Norberg, and Davidson (1998) The location of these separation points is crucial for guiding active flow control strategies. Synthetic jets control flow separation by periodically injecting or extracting fluid into or from the boundary layer, directly altering its momentum distribution.annurevCattafesta; YOU20081349 This enhances the energy within the boundary layer, delays the separation point, and prevents flow separation. The effectiveness of synthetic jets depends largely on parameters such as amplitude, frequency, and application location.SCOTTCOLLIS2004237; Aram2018 Extensive parameter studies are necessary to optimize control performance. The location of the synthetic jet is particularly crucial. Selecting the appropriate location maximizes the jet’s impact on the boundary layer, delays the separation point, and improves control performance. By optimizing the amplitude and frequency of the jet and selecting the best application location, optimal flow control can be achieved, significantly enhancing aerodynamic performance and system stability.
Yan et al., Chen et al. and Yan et al.Yan et al. (2023); Chen et al. (2023); Yan et al. (2024) optimized the position of synthetic jets to maximize control effectiveness. Particularly, two studies have reached starkly different conclusions, prompting further reflection and analysis by researchers. Chen et al.Chen et al. (2023) utilized DRL-based AFC to mitigate vibrations caused by lift on a square cylinder at $Re=100$ and investigated the impact of synthetic jet positioning on control performance at this Reynolds number. Their findings indicate that when synthetic jets are positioned near the leading edge, the drag coefficient is reduced by 7.2% compared to the baseline flow. However, when positioned near the trailing edge, the drag coefficient is reduced by 13.3%. Therefore, at $Re=100$, placing synthetic jets near the trailing corners of the square cylinder is more advantageous for achieving higher drag reduction and lift coefficient reduction. Yan et al.Yan et al. (2023) used DRL and zero-flux jet actuators based on CFD to reduce drag and lift on a square cylinder at $Re=500$, $1,000$, and $2,000$. At $Re=500$, when synthetic jets are positioned near the leading corners, the drag coefficient is reduced by 44.4% compared to the baseline flow. When positioned near the trailing corners, the drag coefficient is reduced by only 1.4%. The results show that placing jets near the leading corners yields better drag reduction than near the trailing corners. Similar conclusions were drawn at $Re=1,000$ and $Re=2,000$. We are curious whether these differing conclusions are due to the different Reynolds numbers or differences in the design of the DRL-based AFC framework.
Given the recent advancements in DRL-based AFC for flow around square cylinders, there is a clear lack of parameterization studies on synthetic jets to reconcile the discrepancies in control effectiveness reported in the literature. Beyond the aforementioned qualitative differences in drag reduction due to synthetic jets positioning at varying Reynolds numbers, we also note that Chen et al.Chen et al. (2023) achieved both drag reduction and complete suppression of vortex shedding in the wake of a cylinder at $Re=100$. However, in Yan et al.Yan et al. (2023), even at Reynolds numbers as high as 500 or more, the control strategies failed to suppress vortex shedding. To determine whether these divergent conclusions arise from differences in Reynolds numbers or variations in DRL-based AFC framework designs, we employ an AFC framework controlled by a SAC agent, testing at consistent Reynolds numbers. This study will conduct a more extensive parametric analysis using two synthetic jet arrangements and four synthetic jet widths to evaluate the sensitivity of these jet parameters to flow control performance. This will help us gain a comprehensive understanding of how changes in jet parameters affect the effectiveness of flow control strategies.
The structure of this paper is as follows: In Sec. II, we detail the numerical simulation methods for flow around a square cylinder, the DRL algorithm, and how the numerical simulation is used as an environment for DRL training to build the DRL-based AFC framework. In Sec. III, we describe the research results in detail. First, we describe the flow patterns of the Baseline and Base flows. Then, we discuss the convergence performance of the DRL algorithm using CFD as the environment, and how the flow field stability around the square cylinder is enhanced during training. Finally, we analyze the effects of synthetic jets positioning and width on flow control performance at $Re=100$ and $Re=500$. We delve into the flow control mechanisms of synthetic jets design near the leading and trailing corners of the square cylinder and explain the physical mechanisms underlying the qualitative conclusions of this study. In Sec. IV, we summarize the main conclusions of this research and discuss the implications of these findings for the field of flow control.